OpenStudy (anonymous):

Solve the logarithm. (1/5)^(x-1)=(1/2)^(3-x)

OpenStudy (dumbcow):

take log of both sides, then put exponent out front --> (x-1)log(1/5) = (3-x)log(1/2)

OpenStudy (dumbcow):

then combine x terms on one side --> xlog(1/5) + xlog(1/2) = 3log(1/2) + log(1/5)

OpenStudy (dumbcow):

Factor out the x, then divide --> x = [3log(1/2) + log(1/5)] /[ log(1/5) + log(1/2)]

OpenStudy (anonymous):

Another person told me the answer was x= ln(40)/ln(10) ? Does that make sense to you??

OpenStudy (dumbcow):

yes just a matter of combining the logs 3log(1/2) + log(1/5) = log(1/2^3 * 1/5) = log(1/40) = -log(40) log(1/2) + log(1/5) = log(1/2*1/5) = log(1/10) = -log(10)

OpenStudy (anonymous):

Okay! Got it. Thank you!

OpenStudy (dumbcow):

yw